# Pre-Training and Personalized Fine-Tuning via Over-the-Air Federated Meta-Learning: Convergence-Generalization Trade-Offs

### v5

@@ VII Conclusion and Discussion
Future work will investigate these trade-offs in scenarios involving more complicated training algorithms such as fully decentralized implementations, asynchronous updates, and fine-tuning of LLMs, as well as large action models (LAMs). For instance, recent advancements in orthogonal time frequency and space (OTFS) modulation potentially improve performance in dynamic environments like linear time-varying (LTV) channels [52]. Another interesting future direction is to optimize transmit power and the number of channel uses to improve on the convergence-generalization trade-offs.
@@
Lemma .2 (Lemma 4.2 in [15])
@@
Lemma .3 (Lemma 4.3 in [15])
@@
Lemma .4 (Lemma 4.4 in [15])
@@
The proof of Theorem IV.1 is based on the perturbed iterate analysis as in [45]. To this end, we define the maintained virtual sequence $\{\hat{\mbox{{$\theta$}}}^{(t)}\}_{t=0,\ldots,T}$ as follows:
@@
The proof of Theorem IV.2 is analogous to that of Theorem IV.1 in principal. It is worth noting that we should carefully design the parameters of $\alpha^{(t)}$ and $\eta^{(t)}$ such that we can arrive at a first-order regression inequality similar to (60). In addition, we modify the memory bound for adaptive learning rate derived in  [53, Lemma 4]. Next, we take a telescopic sum, and use the inequalities $\sum_{t=0}^{T-1}\eta^{(t)}\geq\xi\ln\left((T+a-1)\mathord{\left/\right.\kern-1.2pt}a\right),~\sum_{t=0}^{T-1}(\eta^{(t)})^{2}\leq\xi^{2}\mathord{\left/\right.\kern-1.2pt}(a-1)$ and $\sum_{t=0}^{T-1}(\eta^{(t)})^{3}\leq\xi^{3}\mathord{\left/\right.\kern-1.2pt}2(a-1)^{2},$ which completes the proof.
@@
To validate the generalization capacity of the proposed Air-meta-pFL, we conduct additional experiments on CIFAR10 [54] and Fashion-MNIST [55] data sets with the experimental parameters specified in Table III. The task of each device is to distinguish among $N=2$ different classes, which are pre-assigned to each device from 10 total classes, with $K=32$ shots for each class (see Fig. 3). Other parameters that are not listed here remain the same as in Sec. VI-A. We use a simple CNN network that consists of two convolutional layers with ReLU activation and max pooling, followed by two fully connected layers for 10-class classification.
Fig. 10 and Fig. 10 show the test accuracy versus global rounds $T$ with Air-meta-pFL and other benchmarks on CIFAR10 [R7] and Fashion-MNIST [R8], respectively. This result shows that the proposed Air-meta-pFL achieves comparable performance with vanilla meta-pFL [15], and significantly outperforms FedAvg [8] across different data sets, demonstrating its good generalization capabilities.

### v6

@@ VII Conclusion and Discussion
Future work will investigate these trade-offs in scenarios involving more complicated training algorithms such as fully decentralized implementations, asynchronous updates, and fine-tuning of LLMs, as well as large action models (LAMs). For instance, recent advancements in orthogonal time frequency and space (OTFS) modulation potentially improve performance in dynamic environments like linear time-varying (LTV) channels [rakesh2025over]. Another interesting future direction is to optimize transmit power and the number of channel uses to improve on the convergence-generalization trade-offs.
@@
Lemma .2 (Lemma 4.2 in [fallah2020personalized])
@@
Lemma .3 (Lemma 4.3 in [fallah2020personalized])
@@
Lemma .4 (Lemma 4.4 in [fallah2020personalized])
@@
The proof of Theorem IV.1 is based on the perturbed iterate analysis as in [stich2018sparsified]. To this end, we define the maintained virtual sequence $\{\hat{\mbox{{$\theta$}}}^{(t)}\}_{t=0,\ldots,T}$ as follows:
@@
The proof of Theorem IV.2 is analogous to that of Theorem IV.1 in principal. It is worth noting that we should carefully design the parameters of $\alpha^{(t)}$ and $\eta^{(t)}$ such that we can arrive at a first-order regression inequality similar to (60). In addition, we modify the memory bound for adaptive learning rate derived in  [basu19qsparse, Lemma 4]. Next, we take a telescopic sum, and use the inequalities $\sum_{t=0}^{T-1}\eta^{(t)}\geq\xi\ln\left((T+a-1)\mathord{\left/\right.\kern-1.2pt}a\right),~\sum_{t=0}^{T-1}(\eta^{(t)})^{2}\leq\xi^{2}\mathord{\left/\right.\kern-1.2pt}(a-1)$ and $\sum_{t=0}^{T-1}(\eta^{(t)})^{3}\leq\xi^{3}\mathord{\left/\right.\kern-1.2pt}2(a-1)^{2},$ which completes the proof.
@@
To validate the generalization capacity of the proposed Air-meta-pFL, we conduct additional experiments on CIFAR10 [krizhevsky2009learning] and Fashion-MNIST [xiao2017fashion] data sets with the experimental parameters specified in Table III. The task of each device is to distinguish among $N=2$ different classes, which are pre-assigned to each device from 10 total classes, with $K=32$ shots for each class (see Fig. 3). Other parameters that are not listed here remain the same as in Sec. VI-A. We use a simple CNN network that consists of two convolutional layers with ReLU activation and max pooling, followed by two fully connected layers for 10-class classification.
Fig. 10 and Fig. 10 show the test accuracy versus global rounds $T$ with Air-meta-pFL and other benchmarks on CIFAR10 [R7] and Fashion-MNIST [R8], respectively. This result shows that the proposed Air-meta-pFL achieves comparable performance with vanilla meta-pFL [fallah2020personalized], and significantly outperforms FedAvg [mcmahan2017communication] across different data sets, demonstrating its good generalization capabilities.